# Compositional Generalization Requires More Than Disentangled Representations

### v1

@@ Appendix D Dataset Augmentation
Appendix E Supplementary Experimental Results
E.1 Full Kernel Characterization
@@
In Sec. 3.3, we characterized the model’s OOD generalization behavior from a kernel-based perspective. In Fig. 1(e), we showed the agreement between the similarity (overlap) matrix with the theorized binary factorized kernel, which showed that the model independently “processes” novel input combinations of $x$ and $y$. Here, we provide the full similarity/distance/overlap matrices across the entire generated dataset, ID and OOD. We note that Fig. 1(e) essentially shows a slice of what are shown in Fig. 8. We compare between generated image by models trained with bump-encoding and rate-encoding inputs, and the matrices are constructed based on three metrics: 1) cosine similarity, 2) L2 distance, and 3) pixel overlap. We notice that the kernel matrices of bump-encoded model exhibit stronger “block” structure as compared to those of the rate-encoded model. Here we note that the matrices have interesting diagonal structure from the non-finite width of the Gaussian bumps.
E.2 MNIST Digit Image Rotation Experiments
@@
E.3 Volume Metric Plots and Generated OOD Images
@@
Figure 10: Comparison of generated OOD images for different network depths between networks with bump-based (left) and rate-based (right) input encodings.
@@
E.4 Neural Tuning Curves

### v2

@@ Appendix E Dataset Augmentation
Appendix F Supplementary Experimental Results
F.1 Full Kernel Characterization
@@
In Sec. 3.3, we characterized the model’s OOD generalization behavior from a kernel-based perspective. In Fig. 1(e), we showed the agreement between the similarity (overlap) matrix with the theorized binary factorized kernel, which showed that the model independently “processes” novel input combinations of $x$ and $y$. Here, we provide the full similarity/distance/overlap matrices across the entire generated dataset, ID and OOD. We note that Fig. 1(e) essentially shows a slice of what are shown in Fig. 8. We compare between generated image by models trained with bump-encoding and ramp-encoding inputs, and the matrices are constructed based on three metrics: 1) cosine similarity, 2) L2 distance, and 3) pixel overlap. We notice that the kernel matrices of bump-encoded model exhibit stronger “block” structure as compared to those of the ramp-encoded model. Here we note that the matrices have interesting diagonal structure from the non-finite width of the Gaussian bumps.
F.2 MNIST Digit Image Rotation Experiments
@@
F.3 Volume Metric Plots and Generated OOD Images
@@
Figure 10: Comparison of generated OOD images for different network depths between networks with bump-based (left) and ramp-based (right) input encodings.
@@
F.4 Neural Tuning Curves